Can you help with this question?
A number added to its reciprocal is 5/2. What is the number?

Answer by nerdybill(7384)

A number added to its reciprocal is 5/2. What is the number?
.
Let x = the number
then
1/x = the reciprocal
.
x + 1/x = 5/2
multiplying both sides 2x:
2x^2 + 2 = 5x
2x^2 -5x + 2 = 0
2x^2 -4x -x + 2 = 0
(2x^2 -4x) - (x - 2) = 0
2x(x - 2) - (x - 2) = 0
(x - 2)(2x - 1) = 0
x = {2, 1/2}
.
Answer can be 2 or 1/2